The length of a garden is 20 yards more than its width. The area is 399 square yards. What are the dimensions? This is on a worksheet but my teacher didn't yeah us how to do theis. Help?
Answer by TimothyLamb(4379)

---
L = w + 20
Lw = 399
---
Lw = 399
(w + 20)w = 399
ww + 20w - 399 = 0
---
the above quadratic equation is in standard form, with a=1, b=20, and c=-399
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 20 -399
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the two real roots (i.e. the two x-intercepts), of the quadratic are:
w = 12.3383079
w = -32.3383079
---
negative width doesn't make sense, so use the positive root
---
w = 12.3383079 yrd
L = 32.3383079 yrd
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php